OpenStudy (anonymous):

How to put 2y=12x+7 in General form: ax+by+c=0

OpenStudy (campbell_st):

put all terms on the same side... 12x - 2y + 7 = 0

OpenStudy (anonymous):

-12x+2y-7=0

OpenStudy (anonymous):

does it have a negative up front?

OpenStudy (anonymous):

0 = 12x -2y + 7 aX + bY + Constant = 0